OpenStudy (anonymous):

How do you find the diameter of a circle that has circumference of 81 meters?

jimthompson5910 (jim_thompson5910):

By using the formula C = pi*d where C = circumference of circle pi = 3.14 (approximately) d = diameter of circle

OpenStudy (anonymous):

I don't think I did it right. I got 25,434. What did you get?

jimthompson5910 (jim_thompson5910):

C = pi*d 81 = pi*d ... plug in the given circumference, solve for d 81/pi = d d = 81/pi d = 81/3.14 d = 25.796178343949 That's approximate of course. Use more decimal digits of pi to get a more accurate answer.

OpenStudy (anonymous):

ok.

jimthompson5910 (jim_thompson5910):

not sure how you got 25,434

jimthompson5910 (jim_thompson5910):

maybe you multiplied?